### v1

# TF-Mamba: A Time-Frequency Network for Sound Source Localization

Abstract
Sound source localization (SSL) determines the position of sound sources using multi-channel audio data. It is commonly used to improve speech enhancement and separation. Extracting spatial features is crucial for SSL, especially in challenging acoustic environments. Previous studies performed well based on long short-term memory models. Recently, a novel scalable SSM referred to as Mamba demonstrated notable performance across various sequence-based modalities, including audio and speech. This study introduces the Mamba for SSL tasks. We consider the Mamba-based model to analyze spatial features from speech signals by fusing both time and frequency features, and we develop an SSL system called TF-Mamba. This system integrates time and frequency fusion, with Bidirectional Mamba managing both time-wise and frequency-wise processing. We conduct the experiments on the simulated dataset and the LOCATA dataset. Experiments show that TF-Mamba significantly outperforms other advanced methods on simulated and real-world data.
Index Terms:
DOA Estimation, Sound Source Localization, Mamba, State-space Model
I Introduction
Sound source localization (SSL) [1] aims to automatically identify the origins of sound by analyzing signals captured through a microphone array. The core functionality of SSL systems is to calculate the angles at which sounds reach the microphones. This spatial information is essential for many speech-based applications, including speech separation [2, 3, 4], speech recognition [5], and speech enhancement [6]. Accurate sound localization can assist these applications to improve their performance. As a result, developing more advanced SSL techniques has become a key focus in research, aiming for greater robustness and adaptability in real-world.
Traditional SSL methods estimate spatial features linked to direct-path signal propagation to map features to source locations. Common spatial features include time delay, inter-channel phase/level difference (IPD/ILD) [7], and relative transfer function (RTF) [8]. For example, the steered response power (SRP) based methods, particularly SRP-PHAT [9], which is obtained from the SRP by applying phase transform whitening, have been foundational due to their simplicity and versatility. These spatial features are easy to estimate in ideal conditions but become challenging in real-world scenarios with noise, reverberation, and multiple moving sources. Because the noise and overlapping speech cause distortions, while reverberation masks or colors the signal. Additionally, moving sources create time-varying spatial cues, also leading to a significant decline in localization accuracy.
In recent years, deep learning methods for localization have gained more attention [10, 11, 12] than traditional methods. These methods approach the task as either feature/location regression or location classification. Common architectures for moving SSL include convolutional neural networks (CNNs) [13, 14] and recurrent neural networks (RNNs) [15, 16]. CNNs extract local spatial features, while RNNs capture long-term temporal context. Deep learning models can take input at the signal level (e.g., time-domain signals, spectrograms) or feature level (e.g., inter-channel phase difference, generalized cross-correlation (GCC) [17], spatial spectrum). Motivated by speech enhancement, the FN-SSL [18] framework builds on previous work, recognizes that the direct-path inter-channel phase difference can be estimated in both narrow-band (frequencies) and full-band (time-frames). It uses dedicated LSTM layers to process each band and show advanced performance.
While LSTM-based models like FN-SSL have shown promising results, a new neural state space model (SSM)-based architecture called Mamba [19, 20] has emerged and achieves superior performance to state-of-the-art models across various tasks [21, 22]. In speech processing, there have been attempts to replace transformers and RNNs with Mamba for tasks like speech separation and speech enhancement [23, 24, 25]. Mamba demonstrates a stronger ability to model extremely long-range dependencies compared to LSTM and is notable for its efficient use of computational resources.
In this work, we introduce Mamba to SSL by proposing a novel architecture referred to as TF-Mamba. It is built on the robust FN-SSL framework, which uses LSTM for time and frequency feature fusion. By replacing the LSTM blocks in FN-SSL with bidirectional Mamba (BiMamba) blocks, TF-Mamba aims to improve audio sequence context modeling while maintaining linear complexity with sequence length. We consider both simulated and LOCATA datasets for our studies to evaluate the effectiveness of proposed TF-Mamba for real-world scenarios. To the best of our knowledge, this is the first attempt to use a state-space-based model for SSL.
Figure 1: Architecture of the proposed TF-Mamba network. Each TF-Mamba block includes a temporal Mamba (T-Mamba) and frequency Mamba (F-Mamba) layer, with skip connections to prevent information loss. Bidirectional Mamba (BiMamba) layers capture both past and future information. “S” denotes the SiLU activation and “Linear” indicates linear projection.
II Related Works
II-A Mamba for Speech
Mamba has demonstrated transformer-level performance across various sequence-based modalities, including audio and speech, which are naturally sequential in waveform or spectrogram forms. Early works applied Mamba to single tasks such as speech enhancement [26, 25], speech separation [23, 24], and audio detection and classification [27, 28]. Additionally, some studies explored self-supervised audio transformers trained with masked spectrogram modeling [29]. The most comprehensive studies to date examine Mamba’s applications in speech enhancement, recognition, synthesis, understanding, and summarization. However, effective model design using SSMs for deep SSL remains unexplored.
II-B Deep Sound Source Localization
In recent years, significant progress has been made in SSL using neural networks. Cross3D [30] uses the SRP-PHAT spatial spectrum as input and employs a causal 3D CNN network for tracking moving sound sources. Similarly, SE-ResNet [31] utilizes a squeeze-and-excitation residual network as an encoder and a gated recurrent unit (GRU) network as a decoder. While SELDnet [32] takes frequency-normalized inter-channel phase difference (IPD) concatenated with the magnitude spectrum as input, SALSA-Lite [33], similar to SELDnet, uses frequency-normalized IPD with the magnitude spectrum as input and a ResNet-GRU network. With advanced performance, the FN-SSL [18] also uses direct path IPD as input and leverages dedicated full-band and narrow-band LSTM layers to exploit inter-band dependencies and inter-channel information.
III Proposed TF-Mamba
III-A Mamba
The structured SSMs [19, 20] efficiently handles long-dependent sequences with low computation and memory needs, serving as a substitute for transformers or RNNs. Mamba [34] enhances SSM by introducing an input-dependent selection mechanism for efficient information filtering and a hardware-aware algorithm that scales linearly with sequence length, enabling faster computation. Mamba’s architecture, combining SSM blocks with linear layers, is simpler but achieves state-of-the-art performance across various long-sequence tasks, offering significant computational efficiency during both training and inference.
The core of Mamba is a linear selective SSM. In the equation below, the variables $h_{t}$, $x_{t}$, and $y_{t}$ represent the hidden state, input, and output at time $t$, respectively. The matrices $A$, $B$, and $C$ are learnable parameters corresponding to the state transition, input, and output processes.
$h_{t}=Ah_{t-1}+Bx_{t},\quad y_{t}=Ch_{t}$ (1)
Due to its linear nature, the entire output sequence $y$ of length $\mathcal{L}$ can be expressed as a convolution between the input sequence $x$ of the same length and a kernel $\mathcal{K}$.
$\quad y=x\ast\mathcal{K},{\text{~{}where~{}}}\mathcal{K}=(CB,AB,\ldots,CA^{\mathcal{L}-1}B)$ (2)
Since the matrices $A_{t}$, $B_{t}$, and $C_{t}$ depend on the input $x_{t}$ (i.e., they are selective), this convolution cannot be computed directly but is instead solved using a parallel scan algorithm. A unidirectional Mamba is an SSM positioned between gated linear layers. For most speech-related tasks, bidirectional modeling is preferred because it captures information from both past and future sequences. Therefore, we use the bidirectional Mamba (BiMamba) proposed in [25], where two SSMs and causal convolutions run in parallel: one processes the original sequence, and the other processes the reversed sequence. The outputs from both SSMs are averaged to incorporate information from both directions.
III-B Network Architecture
The structure of the proposed TF-Mamba network is illustrated in Fig. 1. The input to the network consists of the real and imaginary parts of the STFT coefficients, resulting in an input channel count that is twice the number of microphones. The model has three modules as follows:
III-B1 Feature Encoder
The inputs are first processed by a feature encoder, which consists of a dilated DenseNet [35] core flanked by two convolutional layers to enhance its spectral properties. The output from the feature encoder is then transformed through the Time-Frequency Mamba block.
III-B2 Time-Frequency Mamba
The Time-Frequency Mamba (TF-Mamba) block consists of one temporal Mamba layer and one frequency Mamba layer. Fig. 1 illustrates a network with a TF-Mamba block, with additional blocks easily added by repeating the block multiple ($N$) times.
The temporal Mamba layers process time frames independently, with all frames sharing the same network parameters. The input is a sequence along the frequency axis of a single time frame, allowing the temporal Mamba layers to learn inter-frequency dependencies related to spatial and localization cues. These layers do not capture temporal information, which is instead handled by the subsequent frequency Mamba layers.
The frequency Mamba layers process frequencies independently, with all frequencies also sharing parameters. Differing from the temporal Mamba, the input is a sequence along the time axis of a single frequency, where the frequency Mamba layers estimate direct-path localization features. Estimating these features frequency-wise has been extensively studied in conventional methods such as in [18]. The proposed frequency Mamba layers focus on exploiting this inter-channel information. Additionally, since it is time-varying for moving sound sources, the frequency Mamba layers also learn the temporal evolution of the direct-path localization features.
Mamba is useful because it efficiently models long-range dependencies while maintaining computational efficiency. The ability of Mamba to handle both temporal and frequency-specific features in a structured manner makes it ideal for capturing complex relationships in audio data, such as the time-varying nature of localization cues, which are crucial for accurate SSL. However, the temporal Mamba and frequency Mamba layers are designed to focus on their respective types of information. Therefore the pieces of information from the frequency Mamba layer can be lost after passing through a temporal Mamba layer, and vice versa. To prevent this information loss, we add skip connections. As shown in Fig. 1, the output of each temporal Mamba layer is added to the input of the following temporal Mamba layer. Similarly, the output of each frequency Mamba layer is added to the input of the next frequency Mamba layer. Further, each Mamba layer inside BiMamba also has the skip connection.
III-B3 Output Decoder
The output from the TF-Mamba blocks is then passed through an average pooling module to reduce the frame rate and is subsequently fed into a fully connected layer with $\tanh$ activation, which converts the output to the desired dimension. We employ a spatial spectrum prediction module for DOA estimation as in [36]. This system learns complex relationships from the audio data to predict the potential origin of sound at every degree within a half-circle of the linear microphone array. The resulting spatial spectrum is an 181-point map. We train our model by minimizing the mean squared error (MSE) loss between the model’s output spatial spectrum and the ground truth.
IV Experiment Setting
IV-A Datasets
We evaluated the proposed method on both simulated and real-world datasets, using two microphones to localize the direction of arrival (DOA) within a 180-degree azimuth range.
Simulated dataset: Microphone signals are obtained by convolving room impulse responses (RIRs) with speech source signals [12]. The speech signals are randomly selected from the training, development, and test sets of the LibriSpeech corpus [37], and then used for model training, validation, and testing, respectively. RIRs are generated using the gpuRIR [38]. We followed the settings from [18] to obtain the simulated dataset with moving sources. The room reverberation time (RT60) is randomly set between 0.2 and 0.6 seconds, and the room size ranges from $4\times 2\times 2$ m to $10\times 8\times 5$ m. The moving trajectories of speech sources are generated according to  [30], with each source maintaining a fixed height. Two microphones, placed 8 cm apart, are randomly positioned in the room on the same horizontal plane as the sound source. This dataset includes 1,511 audio files, which amount to approximately 4.96 hours of speech data. Based on the RIRs, we generate the mixture data for each original single-channel speech data for every 5 degrees from 0 to 180 degrees, which expands the audio files by 37 times. We then use white, babble, and factory noises from the NOISEX-92 [39] dataset as noise sources and create a diffuse sound field described in [40]. The generated diffuse noise signals are added to the clean signals with a signal-to-noise ratio (SNR) between -10 and 10 dB for noise robustness.
Real-world dataset: We also tested our proposed method on tasks 3 and 5 of the LOCATA dataset [41] following the setting in [18]. The room size was $7.1\times 9.8\times 3$ m, with a reverberation time of 0.55 s. We used microphones 6 and 9 from the DICIT array [42], which have the same configuration as the simulated microphone array. The models trained on the simulated dataset were directly tested on the LOCATA dataset.
IV-B Implementation details and metrics
For training each task with our proposed TF-Mamba model, we use 100 epochs with a 20-epoch early stopping criterion. We utilize the AdamW optimizer with a learning rate of 0.001. Additionally, we employ StepLR, to adjust the learning rate during training. For STFT computations, the frame size is set to 32 milliseconds, and the hop size to 10 milliseconds. The frequency range for the STFT is between 100 Hz and 8000 Hz, with an n_fft value of 512, which determines the length of the windowed signal after zero-padding.
We use two primary metrics: mean absolute error (MAE) and accuracy (ACC). MAE quantifies the average magnitude of prediction errors, disregarding their direction. ACC, on the other hand, measures the percentage of predictions that are exactly correct or within a specified tolerance. We assess accuracy at two tolerance levels: within a $10^{\circ}$ and $15^{\circ}$ error margin. These tolerance levels provide insight into the model’s precision as well as its practical effectiveness.
TABLE I: Ablation studies of TF-Mamba on simulated data. “Expand” denotes the expand factor for the Mamba layer.
Components Params[M] ACC($15^{\circ}$) [%] ACC($10^{\circ}$) [%] MAE []
Block $\times$ 3 (Expand = [2,2,2]) 0.9 96.5 92.6 2.89
Block $\times$ 4 (Expand = [2,2,4,4]) 1.3 98.4 94.9 2.87
Block $\times$ 5 (Expand = [2,2,4,4,8]) 1.8 98.9 96.9 2.52
w/o BiMamba 1.4 97.1 93.6 2.84
w/o Skip 1.8 97.9 95.4 2.65
IV-C Reference baselines
We compared the following five moving SSL methods: Cross3D [30], SELDnet [32], SE-Resnet [31], SALSA-Lite [33], and FN-SSL [18]. It is noted that the SELDnet, SALSA-Lite, and SE-ResNet are models designed for joint sound event detection and localization; hence, we only utilize their localization branches for comparison. All these methods are reproduced and trained on the same dataset as our proposed method. We adjust the time pooling kernel size in each method to achieve the same output frame rate across all models for consistency.
TABLE II: Performance comparison in ACC and MAE on simulated data.
Methods Params Clean -10 dB
[M] ACC($15^{\circ}$) ACC($10^{\circ}$) MAE[] ACC($15^{\circ}$) ACC($10^{\circ}$) MAE[]
SELDnet 0.8 96.6 93.8 3.5 52.8 50.2 16.7
FN-SSL 2.1 98.3 95.9 2.8 71.6 68.4 8.9
Cross3D 5.6 84.8 81.3 6.5 38.8 35.4 26.3
SE-ResNet 10.2 96.7 95.1 3.2 67.7 64.6 10.8
SALSA-Lite 14.0 96.8 95.6 2.9 69.1 65.7 10.3
TF-Mamba 1.8 98.9 96.9 2.5 74.6 72.5 8.2
TABLE III: Performance comparison in ACC and MAE on LOCATA dataset.
Methods Params[M] ACC($15^{\circ}$) ACC($10^{\circ}$) MAE []
SELDnet 0.8 94.1 91.2 5.3
FN-SSL 2.1 96.7 93.3 3.6
Cross3D 5.6 94.8 91.4 5.1
SE-ResNet 10.2 94.0 90.7 6.5
SALSA-Lite 14.0 95.0 92.5 4.6
TF-Mamba 1.8 97.2 94.3 3.2
V Results and Analysis
V-A Ablation: Finding the best configuration for TF-Mamba
The ablation study results for TF-Mamba, shown in Table I, highlight the importance of different components in the network. “Block $\times$ $N$” refers to stacking $N$ number of TF-Mamba blocks. As the number of blocks increases from 3 to 5, accuracy ($10^{\circ}$ error tolerances) improves from 92.6% to 96.9%, showing that deeper networks help the model understand more complex localization information. The “w/o BiMamba” variant, which replaces the bidirectional mamba with unidirectional, reduces accuracy to 93.6% ($10^{\circ}$) and 97.1% ($15^{\circ}$). This shows that processing both past and future information is important for achieving good performance. The “w/o Skip” variant, which excludes skip connections, also leads to a small drop in accuracy and a higher error. This confirms that skip connections are helpful in keeping information flowing through the network. We use Block $\times$ 5 version with BiMamba and skip connections in the following comparison experiments.
V-B Results on simulated data
We report the results in two acoustic conditions of testing in Table II which are “Clean” and “SNR=-10 dB”. The results demonstrate that TF-Mamba outperforms other methods in clean and noisy environments, highlighting its robustness in handling noise and reverberation. With the highest accuracy in $10^{\circ}$ (96.9%), $15^{\circ}$ (98.9%) and lowest MAE (2.5) in clean conditions, TF-Mamba outperforms FN-SSL. Although FN-SSL also analyzes direct-path features, TF-Mamba captures time and frequency dependencies more effectively, leading to better accuracy and robustness in challenging environments. In contrast, methods like SALSA-Lite, which processes noisy IPD with magnitude spectrum, and Cross3D, which uses the SRP-PHAT spatial spectrum, show significant performance drops under noise, underscoring the advantage of TF-Mamba.
V-C Results on real-world data
The results on the LOCATA dataset are presented in Table III, where the models trained on simulated data are directly tested. We observe that TF-Mamba achieves the best performance, with the highest accuracy at both 15° (97.2%) and 10° (94.3%) error tolerances, and the lowest MAE (3.2). This dataset’s relatively good acoustic conditions, which are almost noise-free and have an RT60 of 0.55s, allow all models to perform well. However, TF-Mamba’s superior accuracy and lower error indicate its ability to better capture fine localization details compared to the other methods. Notably, methods like SE-ResNet and SALSA-Lite, which are more complex in terms of parameters, show higher MAE, while simpler models like SELDnet struggle with accuracy and MAE. TF-Mamba’s approach also avoids the over-smoothing problem seen in other models during sudden sound source turns, maintaining strong localization performance in real-world conditions.
VI Conclusion
In this paper, we introduced TF-Mamba, a novel architecture for SSL that leverages the Mamba SSM to fuse time and frequency features effectively. By replacing traditional LSTM blocks with bidirectional Mamba, TF-Mamba enhances the modeling of audio sequence contexts while maintaining computational efficiency. Our experimental results, conducted on both simulated and real-world datasets, demonstrate that TF-Mamba significantly outperforms existing state-of-the-art methods. These findings highlight the potential of SSM like Mamba in advancing SSL tasks, marking the first application of such models in this field.
VII Acknowledgement
The authors would like to thank Yabo Wang et. al. (The authors of FN-SSL [18]) for their valuable suggestions during this work on reproducing the FN-SSL model and other baseline models.